1.The half-life of a radioisotope is found to be 4.55 minutes. If the decay follows first order kinetics, what percentage of isotope will remain after 2.00 hours?

Grade:12th pass

2 Answers

Meghna roy Chowdhury
8 years ago
T1/2 = 4.55min.
 
t1/2= 0.693/k .
hence k= 0.15 min^-1
 
let percentage decomposition be x
t=2 hrs = 120min
 
formula: t=( 2.303 log (100/(100-x)))/k.
 
on substituting known values , we get x = 99.99% 
Denisha
6 years ago
2hrs*60min/1hr*1 half life/4.55 mins=26 half lives
For our total answer we get .5^26 or 1/2^26 which equals 1.5 * 10^-8 %